Telescope

F/35 rebuild

The f/35 secondary chopper electronics were largely rebuilt after the accident at the start of the year. Much of the cabling was replaced due to damage and reconfigured to reduce significantly system noise and susceptibility to EMI and to make troubleshooting easier. New LVDT's for mirror position sensing were added. The voice coil drivers were replaced with spares, and the damaged units were rebuilt to serve as backups. System mechanics were cleaned, maintained and secured in the Waimea shops before final assembly. The chopping system was thoroughly tuned and tested using a dummy mirror.

The repaired and re-coated secondary mirror was returned from Contraves with essentially the same optical figure as when initially procured according to interferograms provided by Contraves. Intense effort by the electronics and mechanics groups had the system ready at the summit about 1 week before the scheduled July IR observing runs. The system has worked essentially without problem since installation.

F/8 Secondary mirror

The F/8 secondary mirror was successfully re-aluminized and put back into service in August. The aluminizing chamber was prepared for the next coating.

Upper end seating

Problems continue with the seating of upper ends on the telescope structure during upper end changes. Clearances between the upper end and the telescope's 'bird's head' locks which are designed to be roughly 13 mm are really on the order of 1-2 mm and can occasionally be zero or less. The result is that the upper end 'hang ups' while handling - a very serious condition which can result in a violent seating of the upper end. The problem is exacerbated by difficulties in accessing the top of the telescope and by the fact that the upper end handling ring completely envelopes the locking system when in the unlocked (problematic) condition.

Inspection of locking segments on the upper end handler revealed several incorrectly sized bolts which resulted in erratic handling ring operation. These were replaced and re-torqued.

Bird's head clearances were investigated with all three top ends on June 5 and 6. Binding in the locking system at this time precluded provision of adequate clearances in the unlocked position. We currently ensure adequate mounting clearances by detailed visual inspection as the top end is put in place. However this lengthy process is not appropriate for the long term.

A short scheduled 3 day shutdown in early September was used for 'bird's head' lock maintenance and investigation. This is the first time since the telescope was assembled that the locks have been disassembled. Custom handling equipment was designed and fabricated to permit removal of the 400 kg locking segments. This equipment worked extremely well. We also purchased a special 3-man 'basket' with dedicated safety harnesses for work around the top of the telescope. Maintenance however is still very crane- intensive.

During this shutdown we were able to service only one of 4 segments in 3 days. Several of 14 wear plates associated with the SE segment showed signs of heavy wear and cracking. Guide bolts were of non- uniform size. Damaged wear plates were replaced. Guide bolts on all 4 segments were inspected and undersized bolts were fitted with precision sleeves. Although this maintenance was clearly overdue, locking systems clearance problems were not much improved. Clearance problems are accentuated by the design of the handling ring's segments which allows the ends of each segment to swing. The swing results in interference as upper ends are seated.

Clearly these systems require a lot more work. We are planning major upgrade/ repair efforts for next summer. In the mean time detailed documentation of as-built dimensions appears to be the only means for resolving inconsistencies between design and operational clearances.

Prime Focus Bonnette

The sources of ongoing intermittent control problem with Z' focus control was identified and solved in June. Old wiring and poor design of its control system continue to make the prime focus bonnette rather unreliable. We are currently pursuing plans to rebuild the bonnette control system in 1997.

The Primary Mirror

Instabilities in the Primary Mirror axial support which have limited use of the telescope to relatively low air- masses have been a severe limitation for IR observers. A detailed review and adjustment of the support systems pneumatic regulators in July largely solved this problem. The piston of one regulator was found to drag slightly. Regulator inlet air pressures were adjusted to establish regulator piston heights according to manufacturer specifications. The telescope can now be used to z = 77° without the primary mirror support system oscillating.

The primary mirror was liquid washed in mid June resulting in a 4.3 % increase of reflectivity and significant reduction in scattered light.

Prime Focus Cage

Dedicated dry air lines and CCD fiber optics cables were run through the cable wraps. We no longer have these systems taped to the outside of the outside cage.

The Telescope Control System

The tacho on an hour angle drive motor failed in mid-semester. Although no significant telescope time was lost, high-current oscillation during telescope slews caused concern until the source of the problem was identified. The faulty motor (tachos are integral to the motor) has been returned to Inland Motors for repair. Several motors - both active and spares - have tacho problems and will be maintained in the coming year. Turn around time is on the order of 4 to 6 months.

An initial thought on the cause of the oscillations was that they might originate with contamination in oil-pad bearing clearances. Accordingly all bearing gaps were thoroughly inspected and found to be well within expected limits.

The spare video image integrator unit (known as Leaky2) that was procured many years ago was tested, documented and installed at the summit as a "hot spare" to the main video integrator used by the TCS autoguider system. Now if the main video integrator (known as Leaky1) fails (for which there have been several scares), the telescope operator can bring Leaky2 on-line with only a few minutes of down time before autoguiding is restored. The TCS4 project will replace these custom video integrators with a commercial image processing module.

TCS autoguider cabling: connections between the video integrator unit (Leaky1) and the data acquisition system (CAMAC) required cleaning and rebuilding of cable connectors.





tmca@cfht.hawaii.edu
Copyright © 1997, Canada-France-Hawaii Telescope Corp. All rights reserved.